0.7 = 3.6
9.8. X
Need to find value of X
My answer is 50.40

Answer by Alan3354(69443)

0.7 = 3.6
9.8. X
Need to find value of X
===========
Enter it like this:
0.7/9.8 = 3.6/x
Cross multiply
0.7x = 9.8*3.6
x = 9.8*3.6/0.7
x = 50.4